OpenStudy (anonymous):

Factor: x2 - 11x + 18 a: (x + 1)(x - 12) b: (x - 3)(x - 6) c: (x - 2)(x - 9) d: prime

OpenStudy (usukidoll):

now this time we need all combinations of 18 1 x 18 = 18 2 x 9 = 18 6 x 3 = 18 so out of all the combinations, which 2 numbers will give me 11 (the numbers are on the left side of the equation)

OpenStudy (usukidoll):

so what's 1 +18 ? 2+9 ? 6+ 3 ?

OpenStudy (anonymous):

it's c

OpenStudy (usukidoll):

can you show me why it's c?

OpenStudy (usukidoll):

I want to make sure you understand the process of factorization

OpenStudy (anonymous):

Find the two factors of 18 that when added together equal 11: -9x-2=18 and -9+-2=-11, so the two factors are -9 and -2. (x-9) and (x-2) set the two equations to 0 and solve: x=9 and x=2.

OpenStudy (usukidoll):

wow that's good (x-9)(x-2) = x^2-2x-9x+18 = x^2-11x+18

OpenStudy (usukidoll):

good job :)